# Trajectory Prediction for Autonomous Driving: Progress, Limitations, and Future Directions

### v1

@@ 6.2.1 Goal Anchor-Based Prediction Methods
To predict plausible goal points as anchors for trajectory generation, models utilize approaches that consider the current context and agent behavior. Once these goals are identified, trajectories are conditioned on them, typically using a path-planning or motion-generation module. Examples include forecasting key destinations, such as parking spots or intersections, and employing goal-conditioned policies to create paths toward these objectives. A key advantage of this approach is its natural alignment with goal-driven tasks, like reaching a specific destination. Goals are often estimated based on static environmental features, including road topology, lane structures, or sidewalks. In road networks, goals can be defined as lane endpoints, intersections, or turn exits, providing a structured basis for trajectory prediction.
Among SOTA data-driven trajectory prediction methods, a set of studies integrate goal-directed principles from planning to infer potential goals and generate goal-conditioned trajectories [539, 81, 532, 540]. By estimating plausible goals, these approaches significantly enhance trajectory prediction accuracy. The BiTraP model [179] utilizes a bidirectional decoder conditioned on predicted goals to reduce accumulated errors. The SGNet framework [535] introduces a stepwise goal estimator that infers and leverages goals at multiple temporal scales. Similarly, PECNet [532] focuses on predicting distant goals to improve the accuracy of long-term trajectory predictions.
@@ 6.2.2 Trajectory Anchor-Based Prediction
Trajectory anchor-based methods focus on predicting potential paths (trajectories) that a vehicle might take to reach its goals [541, 542]. TPNet model [297] generates feasible future trajectories using polynomial curves. Inspired by the sampling-based paradigm in vehicle motion planning [543, 544], PRIME combines planning-based approaches with learning-based prediction by directly searching for reachable paths and generating potential trajectory anchors. This hybrid approach handles complex agent-map interactions while satisfying environmental and kinematic constraints [514], leveraging the strengths of both model-based and learning-based methods for optimal performance in complex scenarios. However, while much focus has been placed on improving point-level prediction accuracy, data-driven frameworks often struggle to enforce constraints during trajectory generation. To address this, Li et al. [545] proposed the Planning-inspired Hierarchical (PiH) framework, which employs a hierarchical lateral and longitudinal decomposition to generate path and goal intentions for agents. This framework directly influences the ego vehicle’s trajectory planning, ensuring alignment with predicted path intentions and promoting safe and efficient navigation.
@@ 6.3 Summary and Discussion
This section reviewed planning-conditioned trajectory prediction methods, focusing on their classifications, methodologies, and contributions to autonomous driving. These methods leverage motion planning principles to predict future trajectories while ensuring constraint compliance and contextual awareness. The reviewed approaches encompass model-based, learning-based, and hybrid techniques, along with anchor-based methods such as goal and trajectory anchors, each presenting distinct advantages and limitations in trajectory prediction.
Model-based approaches employ explicit mathematical representations of agent dynamics and environmental constraints to produce interpretable and physically feasible trajectories. Although reliable and constraint-compliant, these methods often lack adaptability to unstructured or highly dynamic environments. In contrast, learning-based methods excel in capturing complex interactions and multimodal behaviors by employing data-driven techniques, but their reliance on large datasets and inability to explicitly enforce safety constraints remain significant challenges. Anchor-based methods, including goal and trajectory anchors, offer a structured approach to trajectory prediction by aligning with real-world driving objectives. While effective in generating diverse trajectory options, these methods may encounter scalability and computational efficiency issues in dense and highly interactive scenarios.
Future research in planning-conditioned trajectory prediction should address these challenges by adopting hybrid frameworks that combine the robustness and interpretability of model-based methods with the adaptability and scalability of learning-based approaches. For instance, methods like PRIME [514], which integrate dynamic trajectory anchors with data-driven evaluation, exemplify the potential of such hybrid approaches to deliver multimodal, constraint-compliant predictions. Additionally, achieving computational efficiency remains critical for real-time applications, particularly on resource-constrained edge devices. Lightweight architectures and parameter-efficient designs can enhance scalability without compromising prediction accuracy. Planning-conditioned trajectory prediction is a crucial area of research that blends strategic motion planning with predictive accuracy. Overcoming current challenges by leveraging hybrid, interaction-aware, and computationally efficient methodologies will be key to advancing autonomous systems.
@@ 7 Vision-Language Models for Prediction
Vision-Language Models (VLMs) are seen as promising technology that could enable vehicles and systems to deeply understand their surroundings, marking a significant advancement in the era of intelligent and explainable transportation. VLMs [546, 547, 548, 549, 550, 551, 552, 553, 554, 555] break down the barriers between textual and visual information to effectively connect multimodal data. In the context of autonomous driving, VLMs can be roughly categorized into: Multimodal-to-Text (M2T) [556, 557, 558, 559] models, which take an image and text as input and generate text as output; Multimodal-to-Vision (M2V) [560, 561] models, where the input is an image and text pair and the output is an image; and Vision-to-Text (V2T) [562, 563] models, which take an image as input and generate text as output. This section mainly focuses on M2T models in the context of perception and prediction tasks. Readers are encouraged to look into the surveys [564, 565] for further details on other VLM configurations and downstream tasks. Figure 13 shows the general flow of a VLM-based prediction model. The model processes two main inputs: visual and textual data. While the image encoder processes visual inputs, e.g., single-camera images, multi-camera feeds, or LiDAR data, extracting spatial and object-level features; the text encoder interprets contextual information, which can include detailed instructions, scene descriptions, specific object inquiries, navigation commands, etc. The encoded features are then passed to a cross-modal attention mechanism [566, 567, 568], which captures the relationships between visual cues and textual instructions to generate joint representations. This representation is further processed by an autonomous driving task-specific network, e.g., as for the prediction task, by a context understanding layer and decoder. The generated future trajectory considers both visual observations and semantic guidance. By leveraging VLMs for autonomous driving tasks, high-level instructions can be aligned with visual perception, resulting in more flexible and human-like driving responses.
@@
This section further reviews VLM-based models designed for trajectory prediction and traffic scene understanding. Given the limited number of prediction models in the literature, we have included a survey of traffic scene understanding models, considering their potential integration into prediction pipelines and their relevance to readers. Next, we provide a list of available language-enhanced AD datasets with brief descriptions. To conclude the section, we discuss the challenges associated with integrating VLMs into the autonomous driving stack.
@@ 7.1 Trajectory Prediction
Table 18 summarizes reviewed VLM and LLM-based models for trajectory prediction. Yang et al. [569] proposed a two-stage evaluation framework. In the first stage, a VLM model analyzes images from six camera views and a BEV visualization with labelled agents to rank the prediction difficulty of agents. To improve the model’s response accuracy, the VLM is provided with several examples of rankings constructed based on displacement errors from the ground truth. In the second stage, the VLM’s output is used to create a subset of training data based on the ranking, resulting in a highly representative and challenging training dataset that enhances the training efficiency of a conventional prediction model. Moon et al. [570] introduced VisionTrap that leverages human gazes and gestures, vehicle turn signals, and road conditions to enhance motion prediction accuracy. The model comprises four main components: a per-agent state encoder that generates agent-specific context features, a visual semantic encoder that encodes images combined with road segments, a text-driven guidance block that guides learning and understanding of detailed visual semantics, and a trajectory decoder. The study includes an inference-time analysis demonstrating the model’s suitability for real-time predictions.
@@
Reference Name Dataset Description Keysan et al. [568] (2023) - NuScenes The prompt describes the target agent and its environment. Yang et al. [569] (2024) - NuScenes Uses VLMs to create a smaller and representative subset of training data to improve the training efficiency of conventional models. Moon et al. [570] (2024) VisionTrap nuScenes-Text Utilizes human gazes and gestures, vehicle turn signals, and road conditions to enhance motion prediction accuracy. Peng et al. [571] (2024) LC-LLM highD Uses LLM as an explainable lane change prediction model with refined Chain-of-Thought (CoT) reasoning. Zheng et al. [572] (2024) TC-Map LLM WOMD Generates Transportation Context Maps as visual prompts and corresponding text prompts to extract situational insights and pass them to a prediction pipeline.
@@
Reference Name Dataset Description Guo et al. [573] (2024) VLM-Auto Custom simulated dataset Uses Qwen-VL model to analyze visual inputs for identifying environmental conditions and generating control parameters to guide driving behaviour. Gopalkrishnan et al. [574] (2024) EM-VLM4AD DriveLM A lightweight multi-frame model for VQA tasks in autonomous driving. Dona et al. [575] (2024) - Waymo PREPER CITY Evaluates hallucinations in pedestrian detection performed by GPT-4V and LLaVA models. Mo et al. [576] (2024) NexusAD CODA-LM A framework based on InternVL2-26B for understanding driving corner cases and generate perception and driving recommendations. Zhou et al. [577] (2025) ELM nuScenes Ego4D Embodied language model (ELM) to enhance scene comprehension using space-aware pre-training and time-aware token selection. Tian et al. [578] (2024) DriveVLM - VLM for planning, comprising modules for scene description scene analysis, and hierarchical planning. Chen and Lu [579] (2024) - 100 images from BDD100k Uses CogVLM model to handle sensors data and employs a Chain-of-Thought (CoT) method to structure reasoning into scene understanding, prediction, and decision-making.
@@ 7.2 Traffic Scene Understanding
Table 19 summarizes reviewed VLM and LLM-based models for scene understanding. Choudhary et al. [559] introduced Talk2BEV, a VLM designed for BEV maps in autonomous driving. This model handles diverse tasks, including predicting traffic participants’ intentions, visual and spatial reasoning, and decision-making based on visual information. Romero et al. [553] argue that using VLMs to query images via natural language addresses issues of expressivity and the need for multiple single-purpose query models, but it often returns redundant and low-quality results. To improve this, they introduced Zelda, a system that adds discriminator and synonym terms to refine user queries into natural language prompts, enhancing accuracy by filtering out low-quality frames. Mo et al. [576] designed a framework for understanding driving corner cases, built upon the foundation model InternVL2-26B and consisting of four phases. For visual perception, the framework uses DINO for 2D detection and DepthAnything for depth estimation. This detection information, along with the original image, is processed by a scene-aware retrieval module to generate contextually relevant information. The Driving Prompt Optimization module then uses this information to guide the model, suppressing hallucinations and focusing on critical scene aspects, ultimately producing responses for perception and driving recommendations. Zhou et al. [577] introduced the Embodied Language Model (ELM) to enhance traditional VLMs’ scene comprehension, supporting tasks such as object localization, detailed scene description, prediction, and event memorization. ELM incorporates space-aware pre-training for robust spatial localization and time-aware token selection to effectively process long-term video data and retrieve relevant past events. Finally, Tian et al. [578] proposed the DriveVLM system for planning in complex environments, with modules for scene description, scene analysis, and hierarchical planning. The scene description module generates linguistic representations of driving conditions and passes these to the scene analysis module, which examines objects’ motion states and behaviours that may influence vehicle actions.
@@ 7.4 Summary and Discussion
VLMs are large-scale models, and the inference time of the VLM-based prediction pipeline is a primary concern when considering integration and computational demands. From an application standpoint, this may necessitate equipping the vehicle with a fully operational computer if inference is conducted locally. Model quantization [589] is a technique employed to reduce model size, offering faster inference at the cost of some accuracy. Other techniques, such as pruning redundant weights [590] and knowledge distillation [591], serve similar purposes. Further experimentation is needed to explore the trade-off between reasoning capability, model size, and inference speed. Hardware-dependent factors, such as parallelism and processor type (TPU, GPU, or CPU), also play a role in optimizing inference time. If not hosted locally, cloud solutions are a possibility. A set of available VLMs [592, 593] offer cloud hosting capabilities, implying that an AV would need to request information via an Application Programming Interface (API). This reliance can lead to latency or may require a shift in the type of computer added to the vehicle, favouring a communication-efficient setup over a computation-heavy station to minimize delays and ensure data transmission with minimal losses.
Another challenge is prompt “engineering”. The alignment of visual and textual modalities within VLMs is facilitated through prompts, leading to the development of prompt engineering —– the process of refining input prompts to guide the model toward a desired behavior or specific output. Providing context to VLMs, alongside the task they need to perform, has a considerable impact on output. For instance, specifying whether the model acts as a “driver” or a “pedestrian” yields different responses in traffic scene scenarios. Additionally, prompts can include instructions on desired output length and format to ensure that only essential information is generated. While the prompt alone may suffice for specific output formatting, few-shot learning is an alternative approach where an example of the expected output format or length is included in the prompt. More advanced techniques, such as Chain-of-Thought prompting, as described in certain models, guide the model through a logical sequence to enhance response quality. An efficient prompt, with reduced redundancies and a clearly specified output format, results in fewer required tokens and, consequently, faster inference times.
Lastly, the non-deterministic nature of some VLMs [594, 595], meaning they can produce varied outputs for the same image-text pair across different runs, requires specific strategies for effective integration. Techniques such as fine-tuning the temperature parameter to balance the model’s creativity, applying majority voting, or averaging results among the top-$k$ outputs across multiple runs are some of the strategies that can be employed. Balancing consistency and flexibility is essential, as VLMs are particularly valuable for addressing edge cases where conventional pipelines may encounter out-of-distribution scenarios. Forcing the model to consistently generate the same output for identical inputs enhances stability and predictability, but it may also limit the model’s ability to adapt to uncertain situations. Therefore, careful evaluation of generative versus deterministic approaches in the context of trajectory prediction is critical. While VLMs show promise for enhancing long-term trajectory predictions, determining their overall utility and optimal configuration within the autonomous driving stack remains an area for ongoing exploration and refinement.

### v2

@@ 6.2.1 Goal Anchor-Based Prediction Methods
In goal anchor-based forecasting methods, plausible goal points are first predicted based on agent behavior and scene context, then used to guide trajectory generation via path-planning or motion-generation modules. Goals are often estimated based on static environmental features, including road topology, lane structures, or sidewalks. In road networks, they can be defined as lane endpoints, intersections, or turn exits, providing a structured basis for trajectory prediction.
Several SOTA prediction methods incorporate goal-directed planning principles to estimate potential destinations and generate goal-conditioned trajectories [536, 81, 529, 537]. By estimating plausible goals, these approaches significantly enhance trajectory prediction accuracy. The BiTraP model [176] utilizes a bidirectional decoder conditioned on predicted goals to reduce accumulated errors. The SGNet framework [532] introduces a stepwise goal estimator that infers and leverages goals at multiple temporal scales. Similarly, PECNet [529] focuses on predicting distant goals to improve the accuracy of long-term trajectory predictions.
@@ 6.2.2 Trajectory Anchor-Based Prediction
Trajectory anchor-based methods focus on predicting potential paths (trajectories) that a vehicle might take to reach its goals [538, 539]. TPNet model [294] generates feasible future trajectories using polynomial curves. Inspired by the sampling-based paradigm in vehicle motion planning [540, 541], PRIME combines planning-based approaches with learning-based prediction by directly searching for reachable paths and generating potential trajectory anchors. This hybrid approach handles complex agent-map interactions while satisfying environmental and kinematic constraints [511], leveraging the strengths of both model-based and learning-based methods for optimal performance in complex scenarios. Next, to enforce constraints during trajectory generation, Li et al. [542] proposed the planning-inspired hierarchical (PiH) framework, which employs a hierarchical lateral and longitudinal decomposition to generate path and goal intentions for agents.
@@ 6.3 Summary and Discussion
This section reviewed planning-conditioned trajectory prediction methods, focusing on their categorization and methodologies. These methods leverage motion planning principles to predict future trajectories while ensuring constraint compliance and contextual awareness. The reviewed approaches encompass model-based, learning-based, and hybrid techniques, along with anchor-based methods such as goal and trajectory anchors, each presenting distinct advantages and limitations in trajectory prediction.
Model-based approaches employ explicit mathematical representations of agent dynamics and environmental constraints to produce interpretable and physically feasible trajectories. Although reliable and constraint-compliant, these methods often lack adaptability to unstructured or highly dynamic environments. In contrast, learning-based methods excel in capturing complex interactions and multimodal behaviors by employing data-driven techniques, but their reliance on large datasets and inability to explicitly enforce safety constraints remain significant challenges. Anchor-based methods, including goal and trajectory anchors, offer a structured approach to trajectory prediction by aligning with real-world driving objectives. While effective in generating diverse trajectory options, these methods may encounter scalability and computational efficiency issues in dense, crowded scenarios.
Future research in planning-conditioned trajectory prediction should address current limitations by developing hybrid frameworks that combine the robustness and interpretability of model-based methods with the adaptability of learning-based approaches, e.g., PRIME [511]. Another important direction is the anchor generation module, which typically analyzes scene context and/or integrates spatial information to infer potential destinations or paths. However, scene context analysis can be computationally expensive and should be evaluated rigorously in terms of inference time. While map-based anchor generation can be more efficient, it must still be tested for scalability as the number of agents increases, since anchors must be generated per agent. The anchors’ generation module is part of the prediction pipeline and should be evaluated online, rather than relying on pre-generated anchors.
@@ 7 Vision-Language Models for Prediction
Vision-Language Models (VLMs) are seen as promising technology that could enable vehicles and systems to deeply understand their surroundings, marking a significant advancement in the era of intelligent and explainable transportation. VLMs [543, 544, 545, 546, 547, 548, 549, 550, 551, 552] break down the barriers between textual and visual information to effectively connect multimodal data. In the context of autonomous driving, VLMs can be roughly categorized into: Multimodal-to-Text (M2T) [553, 554, 555, 556] models, which take an image and text as input and generate text as output; Multimodal-to-Vision (M2V) [557, 558] models, where the input is an image and text pair and the output is an image; and Vision-to-Text (V2T) [559, 560] models, which take an image as input and generate text as output. This section mainly focuses on M2T models in the context of perception and prediction tasks. Readers are encouraged to look into the surveys [561, 562] for further details on other VLM configurations and downstream tasks. Figure 13 shows the general flow of a VLM-based prediction model. The model processes two main inputs: visual and textual data. While the image encoder processes visual inputs, e.g., single-camera images, multi-camera feeds, or LiDAR-range image, extracting spatial and object-level features; the text encoder interprets contextual information, which can include detailed instructions, scene descriptions, specific object inquiries, navigation commands, etc. The encoded features are then passed to a cross-modal attention mechanism [563, 564, 565], which captures the relationships between visual cues and textual instructions to generate joint representations. This representation is further processed by an autonomous driving task-specific network, e.g., as for the prediction task, by a context understanding layer and decoder. The generated future trajectory considers both visual observations and semantic guidance.
@@
This section reviews VLM-based models designed for trajectory prediction and traffic scene understanding. Given the limited number of prediction models in the literature, we have included a survey of traffic scene understanding models, considering their potential integration into prediction pipelines. The section also provides a list of available language-enhanced AD datasets with brief descriptions and discusses the challenges associated with integrating VLMs into the autonomous driving stack.
@@ 7.1 Trajectory Prediction
Table 18 summarizes reviewed VLM and LLM-based models for trajectory prediction. Yang et al. [566] proposed a two-stage evaluation framework. In the first stage, a VLM model analyzes images from six camera views and a BEV visualization with labeled agents to rank the prediction difficulty. To improve the model’s response accuracy, the VLM is provided with several examples of rankings constructed based on displacement errors from the ground truth. In the second stage, the VLM’s output is used to create a subset of training data based on the ranking, resulting in a highly representative and challenging training dataset that enhances the training efficiency of a conventional prediction model. Moon et al. [567] introduced VisionTrap that leverages human gazes and gestures, vehicle turn signals, and road conditions to enhance motion prediction accuracy. The model comprises four main components: a per-agent state encoder that generates agent-specific context features, a visual semantic encoder that encodes images combined with road segments, a text-driven guidance block that guides learning and understanding of detailed visual semantics, and a trajectory decoder. The study includes an inference-time analysis demonstrating the model’s suitability for real-time predictions.
@@
Reference Name Dataset Description Keysan et al. [565] (2023) - NuScenes The prompt describes the target agent and its environment. Yang et al. [566] (2024) - NuScenes Uses VLM to create a smaller and representative subset of training data to improve the training efficiency of conventional models. Moon et al. [567] (2024) VisionTrap nuScenes-Text Utilizes human gazes and gestures, vehicle turn signals, and road conditions to enhance motion prediction accuracy. Peng et al. [568] (2024) LC-LLM highD Uses LLM as an explainable lane change prediction model with refined Chain-of-Thought (CoT) reasoning. Zheng et al. [569] (2024) TC-Map LLM WOMD Generates Transportation Context Maps as visual prompts and corresponding text prompts to extract situational insights and pass them to a prediction pipeline.
@@
Reference Name Dataset Description Guo et al. [570] (2024) VLM-Auto Custom simulated dataset Uses Qwen-VL model to analyze visual inputs for identifying environmental conditions and generating control parameters to guide driving behaviour. Gopalkrishnan et al. [571] (2024) EM-VLM4AD DriveLM A lightweight multi-frame model for VQA tasks in autonomous driving. Dona et al. [572] (2024) - Waymo PREPER CITY Evaluates hallucinations in pedestrian detection performed by GPT-4V and LLaVA models. Mo et al. [573] (2024) NexusAD CODA-LM A framework based on InternVL2-26B for understanding driving corner cases and generate perception and driving recommendations. Zhou et al. [574] (2025) ELM nuScenes Ego4D Embodied language model (ELM) to enhance scene comprehension using space-aware pre-training and time-aware token selection. Tian et al. [575] (2024) DriveVLM - VLM for planning, comprising modules for scene description scene analysis, and hierarchical planning. Chen and Lu [576] (2024) - 100 images from BDD100k Utilizes the CogVLM model to process sensor data and employs a Chain-of-Thought (CoT) approach to structure reasoning across scene understanding, prediction, and decision-making.
@@ 7.2 Traffic Scene Understanding
Table 19 summarizes reviewed VLM and LLM-based models for scene understanding. Choudhary et al. [556] introduced Talk2BEV — a VLM designed to handle diverse tasks, including predicting traffic participants’ intentions, visual and spatial reasoning, and decision-making based on visual information. Romero et al. [550] argue that using VLMs to query images via natural language addresses issues of expressivity and the need for multiple single-purpose query models, but it often returns redundant and low-quality results. To improve this, they introduced Zelda, a system that adds discriminator and synonym terms to refine user queries into natural language prompts, enhancing accuracy by filtering out low-quality frames. Mo et al. [573] designed a framework for understanding driving corner cases, built upon the foundation model InternVL2-26B and consisting of four phases. For visual perception, the framework uses DINO for 2D detection and DepthAnything for depth estimation. This detection information, along with the original image, is processed by a scene-aware retrieval module to generate contextually relevant information. The Driving Prompt Optimization module then uses this information to guide the model, suppressing hallucinations and focusing on critical scene aspects, ultimately producing responses for perception and driving recommendations. Zhou et al. [574] introduced the Embodied Language Model (ELM) to enhance traditional VLMs’ scene comprehension, supporting tasks such as object localization, detailed scene description, prediction, and event memorization. ELM incorporates space-aware pre-training for robust spatial localization and time-aware token selection to effectively process long-term video data and retrieve relevant past events. Finally, Tian et al. [575] proposed the DriveVLM system for planning in complex environments, with modules for scene description, scene analysis, and hierarchical planning. The scene description module generates linguistic representations of driving conditions and passes these to the scene analysis module, which examines objects’ motion states and behaviours that may influence vehicle actions.
@@ 7.4 Summary and Discussion
VLMs are large-scale models, and the inference time of the VLM-based prediction pipeline is a primary concern when considering integration and computational demands. From an application standpoint, this may necessitate upgrading the vehicle’s onboard compute hardware if inference is conducted locally. Model quantization [586] is a technique employed to reduce model size, offering faster inference at the cost of some accuracy. Other techniques, such as pruning redundant weights [587] and knowledge distillation [588], serve similar purposes. Further experimentation is needed to explore the trade-off between reasoning capability, model size, and inference speed. Hardware-dependent factors, such as parallelism and processor type (TPU, GPU, or CPU), also play a role in optimizing inference time. If not hosted locally, cloud solutions are a possibility. A set of available VLMs [589, 590] offer cloud hosting capabilities, implying that an AV would need to request information via an Application Programming Interface (API). This reliance can lead to latency or may require a shift in the type of hardware used onboard, favouring a communication-efficient setup over a computation-heavy station to minimize delays and ensure data transmission with minimal losses.
Another challenge is prompt engineering. It refers to the process of crafting input prompts that guide the model toward specific behaviors or outputs, especially by aligning visual and textual inputs. Providing relevant context within the prompt significantly affects the model’s response. For example, including scene-specific cues such as “heavy traffic ahead” or “intersection with pedestrians” can influence the model’s understanding and downstream reasoning. Prompts can also include instructions on output length or format, helping ensure concise and relevant generation. In some cases, few-shot prompting is used instead, where an example of the expected output is embedded within the prompt. Additionally, techniques such as Chain-of-Thought (CoT) prompting structure the model’s reasoning into intermediate steps, enhancing both interpretability and output quality. An efficient prompt design not only improves semantic alignment but also reduces token usage, contributing to lower inference time.
Lastly, the non-deterministic nature of some VLMs [591, 592], where the same image-text input may yield different outputs across runs, presents integration challenges in safety-critical systems. Strategies such as adjusting the temperature parameter to control output variability, applying majority voting, or averaging across the top-$k$ outputs can help manage this variability. Striking the right balance between consistency and flexibility is essential: enforcing deterministic outputs enhances system stability, while preserving generative flexibility allows the model to better handle edge cases and out-of-distribution scenarios. A rigid deterministic setup may limit adaptability, especially in complex environments. Consequently, the choice between generative and deterministic use of VLMs must be carefully evaluated in the context of trajectory prediction. Although VLMs show promise for enhancing situational understanding and predictive reasoning, their optimal configuration within the autonomous driving stack remains an on-going research investigation.